| Name | Email | Date | Comment
|
| ANS
| anolla@naic.edu
| 5/18/2000
| The High Pass filters were removed from system , they want to observe at lower
frequencies.
|
| ANS
| anolla@naic.edu
| 5/19/2000
| The receiver warmed up during the night time ,went up to checked it.We do not
find anything wrong with the cryo system ,so we purged and cooled it down.Return to
normal.
|
| ANS
| anolla@naic.edu
| 8/4/2000
| The crosshead was replaced due to contamination .It was ED 019 , now is ED
005. Also the R.F. amps & circulators were replaced due to a bad vswr problem. Now
amp 9 is on chan A and amp 8 in on B.
|
| ANS
| anolla@naic.edu
| 8/4/2000
| System was calibrated today , we use the HP D.C. 778 sn 1144A04065 , CU
E107169.
|
| ANS
| anolla@naic.edu
| 8/9/2000
| Dewar's orignal amplifiers were installed back , the circulators are the
replacement ones.Chan. A was adjusted using the NAR mode in order to improve it's
Tsys.Tsys for chan.A is 34 K deg and for chan. B is 29 K deg.
|
| ANS
| anolla@naic.edu
| 8/29/2000
| A new vacpump was installed to the Dewar.It's on the Dewar but not connected
to it's power supply.It's a Duniway vacpump model VA-8-DD/0.
|
| ANS
| anolla@naic.edu
| 1/22/2001
| There is a cryo leak at line 9 in the rotary floor , is about 12 pounds / day.
|
| ANS
| anolla@naic.edu
| 3/12/2001
| Dewar been cooled down due to a cryo comp failure.Cryo comp was off due to a
power glitch on 3/10/2001.
|
| ANS
| anolla@naic.edu
| 4/11/2001
| New NLB filter bank installed , PP checking for RFI.
|
| ANS
| anolla@naic.edu
| 7/27/2001
| Cryo comp ED 012 replaced with ED 010. ED 012 taken down for maint.
|
| ANS
| anolla@naic.edu
| 8/21/2001
| The cryo comp was tripped off yesterday due to a power glitch.
|
| ANS
| anolla@naic.edu
| 11/12/2001
| Cryo comp ED 010 was replace with cryo comp ED 009.
|
| ANS
| anolla@naic.edu
| 11/19/2001
| Cryo comp ED 009 was replaced with cryo comp ED 010, there is not a remote
reset sw on these compressor.
|
| ANS
| anolla@naic.edu
| 11/26/2001
| System was warmed. JR & IC went up to restart it. It was due to power dip on
Sunday.
|
| ANS
| anolla@naic.edu
| 2/14/2002
| System shut off in order to replace cry pump. 430 kHz Dual beam was running.
|
| ANS
| anolla@naic.edu
| 2/15/2002
| Cryopump replaced, was ED 005 and now is ED 020.
|
| ANS
| anolla@naic.edu
| 2/18/2002
| System got hot due to failure of the cryocomp (it's fan quit working overheating
the compressor) So, WG decided to hook to other cryocomp (hooked to the XB, NSB
recvrs comp, ED 009).
|
| ANS
| anolla@naic.edu
| 2/28/2002
| Due to a kinked supply line, (line 7, it was kinked by the cryocomp ED 010 when
we were installing it as spare cryocomp) the emergency cryoline hook up was modified,
it was performed at the cryocomp shack first but later was hooked at the rotary floor
instead.
|
| ANS
| anolla@naic.edu
| 5/13/2002
| Due to power dip cryocomp failed , have to pump it .Normal again.
|
| ANS
| anolla@naic.edu
| 5/15/2002
| installed 1280 MHz waveguide Hi-pass filter after cryo amp and before the 2 post amps.
Checked for ripple at output of postamps, before and after the isolator. Small 1.5 dB
ripple seen in 1200 – 1350 region, should be OK. Tsys looks OK. (29K/30K)
|
| ANS
| anolla@naic.edu
| 5/17/2002
| System was hot due that cryocomp quit because a power dips. Was cooled.
|
| ANS
| anolla@naic.edu
| 5/31/2002
| There is a new filter bank installed with this receiver. There is a new 3488 HP
controller in RF cab 2 in order to control these filter bank. Also we split the power for the
filter bank to share with the one on WLB filter bank at the RF cabinet.
|
| ANS
| anolla@naic.edu
| 7/31/2002
| On 6/21/2002 the cryocomp 009 failed (cooling fan stop working) , it was
replaced by cryocomp 011.
|
| LSW | lwray@naic.edu | Friday, August 16, 2002 15:54:07 | Power glitch caused 3-receiver compressor (LBN, SBN, XB) to warm up around 10am. Detected problem about 1.5 hours later. Rx warmed up past 200K, was cooled by 4:30pm.
|
| ans | anolla@naic.edu | Wednesday, August 21, 2002 13:48:58 | Feedhorn (Dewar) lowered by 3.84" in order to adjust the focal point.New mounting brackets installed.Old brackets are at the Dome.
|
| ans | anolla@naic.edu | Monday, September 9, 2002 15:56:15 | The cryo compressor was shut off by a power dip on 9/8/2002 at 08:30, it was not noticed after several hours .WG have to come to cool it down.
|
| LSW | lwray@naic.edu | Monday, November 4, 2002 15:03:04 | +12V power supply that runs the heater, cal driver, and other things was replaced today. It was old and rusty and possibly contributing to bad baselines (but not proven to be so).
|
| LSW | lwray@naic.edu | Wednesday, November 20, 2002 17:05:50 | Nov 16, 10PM: observer noticed that CH B had no power. Intermittent problem. Probably due to fault in Ch B cryoamp's stage 2 wiring. analyzing problem and possible fixes.
|
| ans | anolla@naic edu. | Monday, January 13, 2003 08:42:37 | Dewar temp up high , 34º. Need cryopump swap.
|
| ans | anolla@naic edu. | Tuesday, January 14, 2003 14:26:04 | Cryopump replaced (was ED 020 . is now ED 013), been cooled.
|
| ans | anolla@naic edu. | Wednesday, February 26, 2003 11:28:56 | Dewar warming up. Cryopump needs replacement.
|
| ans | anolla@naic edu. | Friday, February 28, 2003 09:05:59 | The Dewar cryo temp was reading hi , but we checked the vp meter upstairs the cryo temp it was 18ºk using the vp meter.
|
| lsw | lwray@naic.edu | Monday, March 3, 2003 16:27:47 | Changed attenuators after post amp from: 0/0 dB (A/B) to 4/0 dB, rebalancing levels after upstairs RF switch change.
|
| lsw | lwray@naic.edu | Wednesday, March 26, 2003 16:26:04 | Added XB to the compressor (011), now has SBN/LBN/XB.
|
| lsw | lwray@naic.edu | Wednesday, March 26, 2003 16:27:37 | Cheo changed the dewar monitoring circuit today. Now it reads 8 to 20K, but shows oscillation in temperature which may be real. Still have to add 70K and OMT monitoring capability.
|
| ans | anolla@naic.edu | Tuesday, April 1, 2003 08:30:04 | Cryopump replaced (was ED 025 , is ED 007).System was cooling normally , but Ganesh found that it was increasing in temp again yesterday at 16:35. This morning was reading 39ºk. Going up to check.
|
| ans | anolla@naic.edu | Tuesday, April 1, 2003 14:57:05 | System cooled , VPG was on 1 , but temp indicator was reading high ( 39ºk).Problems with temp indicator Cheo working on it.
|
| ans | anolla@naic.edu | Thursday, April 3, 2003 11:43:14 | The past cryo problems with this receiver pinpoints to faulty 16º stage cap seals.We used this time a CTI 16º stage cap seal and it cooled down to it's normal 16º stage temp , +/- 15 ºk.
We still have problems with it's remote temp indicator, but Cheo is working on it.
|
| ans | anolla@naic.edu | Thursday, April 3, 2003 16:24:33 | Phellip Perillat adjusted the cryo remote setting reference for the LBN Dewar reading. So rigth now is reading 13ºk that is it's reading upstairs.
|
| ans | anolla@naic.edu | Monday, April 7, 2003 14:17:51 | JRV installed on 4/3/2003 the ckt to read both cryotemp stages downstairs.
|
| ans | anolla@naic.edu | Tuesday, April 22, 2003 08:51:18 | Cryo system was hot due to faulty cryopump,it got stock with one of the holding screw of the displacers.Was replaced with an spare one. It cooled down normally. system back to service.
|
| ans | anolla@naic.edu | Wednesday, July 23, 2003 09:02:05 | The system warmed up due to power dip.Operator try to restart the cryocomp but didn,t respond.Found later that restart ckt relay was bad.
|
| ans | anolla@naic.edu | Thursday, October 2, 2003 10:07:14 | The system warmed up due to bad cryopump driver , capacitors failed again.Using old driver and is ok.
|
| ans | anolla@naic.edu | Monday, October 27, 2003 13:19:48 | Shutter removed from horn after dual beam run.
|
| ans | anolla@naic.edu | Monday, November 10, 2003 13:14:35 | System warmed up on 11/8/2003 due to cryocomp failure .JRV came on Saturday to replace cryocomp ED 011 by cryocomp ED 010 (SBH recevr one ).
|
| ans | anolla@naic.edu | Tuesday, November 18, 2003 11:56:11 | System returned to new trio cryocomp ( ED 008), but it's efficiency drop down a bit.
|
| ans | anolla@naic.edu | Wednesday, January 7, 2004 15:02:18 | The NLB receiver removed from operation. The XB will be at that location.
|
| lsw | lwray@naic.edu | Friday, January 16, 2004 10:45:16 | LBN receiver moved to lab. Decomissioned. :(
|